• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name R. Michaels

    Log entry time 00:10:26 on April24,2010

    Entry number 319076

    keyword=putting in thick C12

    23:45 .... Putting in thick C12 target.
    Will run like this until about afternoon sometime.

    Plan is to take two slugs tonight at 50 uA, then
    try to increase the current during the day.

    Watch the temperatures ! If they go too high, reduce the
    beam current.

    The HV's listed in the opmanual HV for C12 which were
    lower than for Pb208. This is wrong. We will adjust
    HV to put the detectors at ~0.65 of the max ADC.

    23:51 .... Ask for 17x7 raster (we were running with bigger
    raster since about 7 pm).

    Bumping up HV

    The mapping of "upper Quartz" (name) to HV channel is from the
    opmanual. The mapping of the HV channel to "det1" (name) is
    from observational reality as we change them.

    Right HRS

    upper Quartz .... chan 4 ... is det2 .... increase from 810 to 950V
    get 150K

    lower Quartz .... chan 6 .... is det1 .... 880 -> 1000. get 150K

    ------------------------
    Left HRS

    lower Quartz .... chan 2 .... is det4 .... 750 -> 865. get 158K

    upper Quartz .... chan 4 .... is det3 .... 850 -> 975. get 146K

    Run 3084 was the run taken while we adjusted HV. Call it junk.

    Run 3085 is the first production run with thick C12.